# Decentralised collaborative action: cryptoeconomics in space

### v1

@@ Abstract
Blockchains and peer-to-peer systems are part of a trend towards computer systems that are radically decentralised, by which we mean that they 1) run across many participants, 2) without central control, and 3) are such that qualities 1 and 2 are essential to the system’s intended use-cases.
We propose a notion of topological space, which we call a semitopology, to help us mathematically model such systems. We treat participants as points in a space, which are organised into actionable coalitions. An actionable coalition is any set of participants who collectively have the resources to collaborate (if they choose) to progress according to the system’s rules, without involving any other participants in the system.
@@
It turns out that much useful information about the system can be obtained just by viewing it as a semitopology and studying its actionable coalitions. For example: we will prove a mathematical sense in which if every actionable coalition of some point $p$ has nonempty intersection with every actionable coalition of another point $q$ — note that this is the negation of the famous Hausdorff separation property from topology — then $p$ and $q$ must remain in agreement. The true power of this observation is that, because it depends only on the semitopological structure, it holds for any possible concrete algorithm.
This is of considerable practical interest, because remaining in agreement is a key correctness property in many distributed systems. For example in blockchain, participants disagreeing is called forking, and blockchain designers try hard to avoid it.
@@ 1 What is ‘cryptoeconomics’?
It began with Napster (a peer-to-peer filesharing system) in June 1999, which demonstrated how music media could be disseminated independently of (centralised) media companies. Then Bitcoin (January 2009) and Ethereum (July 2015) changed everything by showing that money and even contracts could be mediated (albeit imperfectly) independently of banks.
These systems — of which Napster, Bitcoin, and Ethereum are perhaps the best known, but certainly not the only examples — would be meaningless and make no sense if they were implemented in a centralised manner, in much the same way that a pocket calculator or desktop computer would make no sense if it were a mainframe. The decentralisation is not a quality of a thing: it is the point of that thing.
@@ 2.1 The definition
Nature is naturally permissionless (living things do not need permission to be born or die). National voting systems are permissioned (because citizens require certification from the government to be allowed to vote). Peer-to-peer systems (including filesharing and blockchain protocols) are very often, though not always, unpermissioned.
@@
Mathematically speaking, it can be useful to admit general models, including both good and bad ones, so that we can formalise their good and bad behaviour555…which will vary by application; e.g. sometimes all participants should play by the same rules, but in the case of a blockchain bridge we specifically want to admit different rules. and express conditions to include or exclude it.
@@ 2.2 Some informal examples from real life
An actionable coalition is any set containing equal numbers of male leads and female followers.
@@
With reference to our couples dance example: an example of an actionable coalition that is not minimal is a set containing two male leads and two female followers. There are two ways for the participants to pair off to collaborate (i.e. dance).
@@
To get a flavour of our mathematical results, consider a fundamental problem in any decentralised system: consensus; i.e. the problem of ensuring that participants remain in agreement, for some suitable sense of ‘agree’. To take a simple example from blockchain: if we reach a situation where half of the nodes say that we have paid for a service, and the other half say that we have not — then everyone has a problem, because the system has become incoherent and it is not clear how the system as a whole can restore coherence and progress.999coherent (adj.) 1550s, “harmonious;” 1570s, “sticking together,” also “connected, consistent” (of speech, thought, etc.), from French cohérent (16c.), from Latin cohaerentem (nominative cohaerens), present participle of cohaerere “cohere,” from assimilated form of com “together” (see co-) + haerere “to adhere, stick” (etymologyonline: https://www.etymonline.com/word/coherent). This phenomenon is called forking, and blockchain designers really want to avoid it!
@@ 3.1 The definition
The reader will recognise a semitopology as being like a topology [Eng89, Wil70], but without the condition that the intersection of two open sets necessarily be an open set. This reflects the fact that the intersection of two actionable coalitions need not itself be an actionable coalition.
@@
The reader may recall the Hausdorff separation property, typical in topology, that any two distinct points should have disjoint open neighbourhoods. As we shall see from Theorem 3.4, for the study of consensus we are particularly interested in semitopologies with antiseparation properties, of which being intertwined is a canonical such property; being intertwined is the precise negation of being Hausdorff separated.
@@ 4.1 Overview
In current ongoing work, we are using these logics used to declaratively specify, reason about, and debug consensus protocols.
@@ 4.2 Future work
Can logics based on semitopologies be used to accelerate development, and increase confidence in, distributed algorithms, by giving new declarative descriptions of consensus algorithms? This is current work.

### v2

@@ Abstract
Blockchains and peer-to-peer systems are part of a trend towards computer systems based on decentralised collaborative action, by which we mean that they 1) run across many participants, 2) without central control, and 3) are such that qualities 1 and 2 are essential to the system’s intended use cases.
We propose a notion of topological space, which we call a semitopology, to help us mathematically model such systems. We treat participants as points in a space, which are organised into actionable coalitions. An actionable coalition is any set of participants who collectively have the resources to collaborate (if they choose) to progress according to the system’s rules, independently of the rest of the system.
@@
It turns out that much useful information about the system can be obtained just by viewing it as a semitopology and studying its actionable coalitions. For example: we will prove a mathematical sense in which if every actionable coalition of some point $p$ has nonempty intersection with every actionable coalition of another point $q$ — note that this is the negation of the famous Hausdorff separation property from topology — then $p$ and $q$ must remain in agreement. Remarkably, since this observation depends only on the semitopological structure, it holds for any possible concrete algorithm.
This matters because remaining in agreement is a key correctness property in many distributed systems. For example in blockchain, participants disagreeing is called forking, and blockchain designers try hard to avoid it.
@@ 1 What is ‘cryptoeconomics’?
It began with Napster (a peer-to-peer filesharing system) in June 1999, which demonstrated how music media could be disseminated independently of (centralised) media companies [Dav16]. [^2] The article itself illustrates some of the compromises involved in creating and disseminating knowledge. The article’s author is an Associate Professor at a UK University. The publisher’s version is behind a paywall (since the publisher makes money from publishing); an author’s preprint is made freely online by his University (which makes money by employing the author to educate students); and the author most likely wrote the article in-between teaching obligations, for zero marginal cost to his employer (i.e. ‘for free’). To be fair: the publisher’s version looks nicer than the author’s preprint, and the publisher’s website makes the work easy to find; the University survives and its students get taught by a well-informed professor; the author’s preprint is accessible to any reader who can dig it out; and the author enjoyed writing the article. In this sense, all parties — the author, the publisher, the students, the university, the article’s readership, and science itself — benefit from the shambling compromise that is academic publishing, though they might also all complain about the division of that benefit. Then Bitcoin (January 2009) and Ethereum (July 2015) changed everything by showing that money and even contracts could be mediated (albeit imperfectly) independently of banks.
These systems — of which Napster, Bitcoin, and Ethereum are perhaps the best known, but certainly not the only examples — would be meaningless and make no sense if they were implemented in a centralised manner, in much the same way that a pocket calculator or desktop computer would make no sense if it were a mainframe. The decentralisation is not a quality of the thing: it is the point of that thing.
@@ 2.1 The definition
Nature is naturally permissionless (living things do not need permission to be born or die). National voting systems are permissioned (because citizens require government certification to be allowed to vote). Peer-to-peer systems (including filesharing and blockchain protocols) are often, though not always, unpermissioned.
@@
Mathematically speaking, it can be useful to admit general models, including both good and bad ones, so that we can formalise their good and bad behaviour666…which will vary by application; e.g. sometimes all participants should play by the same rules, but in the case of a blockchain bridge we specifically want to admit different rules. and express conditions to include or exclude it.
@@ 2.2 Some informal examples from real life
An actionable coalition is any set containing equal numbers of leads and followers.
@@
With reference to our couples dance example: an example of an actionable coalition that is not minimal is a set containing two leads and two followers. There are two ways for the participants to pair off to collaborate (i.e. dance).
@@
To get a flavour of our mathematical results, consider a fundamental problem in any decentralised system: consensus; i.e. the problem of ensuring that participants remain in agreement, for some suitable sense of ‘agree’. To take a simple example from blockchain: if we reach a situation where half of the nodes say that we have paid for a service, and the other half say that we have not — then everyone has a problem, because the system has become incoherent and it is not clear how the system as a whole can restore coherence and progress.101010coherent (adj.) 1550s, “harmonious;” 1570s, “sticking together,” also “connected, consistent” (of speech, thought, etc.), from French cohérent (16c.), from Latin cohaerentem (nominative cohaerens), present participle of cohaerere “cohere,” from assimilated form of com “together” (see co-) + haerere “to adhere, stick” (etymologyonline: https://www.etymonline.com/word/coherent). This phenomenon is called forking, and blockchain designers really want to avoid it!
@@ 3.1 The definition
We recognise a semitopology as being like a topology [Eng89, Wil70], but without the condition that the intersection of two open sets necessarily be an open set. This reflects the fact that the intersection of two actionable coalitions need not itself be an actionable coalition.
@@
This recalls the Hausdorff separation property, typical in topology, that any two distinct points should have a disjoint pair of open neighbourhoods. As we shall see from Theorem 3.4, for the study of consensus we are particularly interested in semitopologies with antiseparation properties, of which being intertwined is a canonical such property; and it is the precise negation of being Hausdorff separated.
@@ 4.1 Overview
In ongoing work, we are using these logics used to formally specify, reason about, and debug consensus protocols.
@@ 4.2 Future work
Can logics based on semitopologies be used to accelerate development, and increase confidence in, distributed algorithms, by giving new declarative descriptions of consensus algorithms?
This is current work, and so far it has proven useful: most recently we applied a semitopological modal logic to axiomatise the Paxos consensus algorithm [GZ25], and in ongoing research we have used a more advanced version of the logic to formally specify and identify errors in a proposed industrial protocol (Heterogeneous Paxos [SWvRM20, SWvRM21]). We are now using the same techniques to help design its replacement.